write an equation that is parallel to the graph of the line that passes though the points (-8,2) and (-3,3)?
Answer by stanbon(75887)

write an equation that is parallel to the graph of the line that passes though the points (-8,2) and (-3,3)?
----
slope = (3-2)/(-3--8) = 1/5
----
Equation:
y = (1/5)x
=========================
Cheers,
Stan H.
=============